OpenStudy (anonymous):

solve by substituting a variable to simplify the equation: x^4+2x^2=8 How do i get the answer; + or - the square root of 2 and + or - 2i

OpenStudy (slaaibak):

use the substitution k = x^2

OpenStudy (anonymous):

\[x^4+2x^2-8=0\] \[u=x^2\] \[u^2+2u-8=0\\]

OpenStudy (anonymous):

how do i get the answer; + or - the square root of 2 and + or - 2i